# Homogenized C. elegans Neural Activity and Connectivity Data

### v3

@@ Abstract
There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C. elegans), as this small model system provides a path to bridge the gap between nervous system structure (connectivity) and function (physiology). However, existing physiology datasets, whether involving passive recording or stimulation, are in distinct formats, and connectome datasets require preprocessing before analysis can commence. Here we compile and homogenize datasets of neural activity and connectivity. Our neural activity dataset is derived from 11 C. elegans neuroimaging experiments, while our connectivity dataset is compiled from 9 connectome annotations based on 3 primary electron microscopy studies and 1 signal propagation study. Physiology datasets, collected under varying protocols, measure calcium fluorescence in labeled subsets of the worm’s 300 neurons. Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling rate, yielding recordings from approximately 900 worms and 250 uniquely labeled neurons. The connectome datasets, collected from electron microscopy reconstructions, represent the entire nervous system as a graph of connections. Our collection is accessible on HuggingFace, facilitating analysis of the structure-function relationship in biology using modern neural network architectures and enabling cross-lab and cross-animal comparisons.
@@ 1 Introduction
To address these challenges, we have created two integrated datasets. The first aggregates neural dynamics from 11 calcium imaging studies, and the second compiles connectome data from three electron microscopy wiring studies [35, 3, 36] and one functional signal propagation study [14]. Together, these datasets combine structure and function, offering an invaluable resource for computational modelers aiming to construct biologically grounded neural network models.
@@ 2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions, files and links) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data into standardized formats.
@@
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The pickle_neural_data() function extracts and processes raw calcium fluorescence data from each of our 11 neural activity source datasets, standardizing them into neural activity matrices that are saved as compressed files for efficient storage and retrieval. The preprocess_connectome() function processes raw connectome data from source files into graph tensors, which are formatted for compatibility with graph neural network (GNN) frameworks such as PyTorch Geometric [21]. This preprocessing step facilitates downstream tasks like connectivity-informed modeling of neural activity. Although there are 10 distinct connectome source files, as detailed in Table 5, all are derived from 4 primary publications [35, 3, 36, 14], ensuring comprehensive coverage of C. elegans neural connectivity data.
Therefore, our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
Figure 1: Overview of Neural Activity Datasets. (A) Distribution of the number of worms in each experimental dataset. (B) Number of recorded neurons per worm compared to the total neuronal population of C. elegans. (C) Total duration of recorded neural activity for each dataset. (D) Average duration of recorded activity per worm, with one hour of calcium imaging as a benchmark. (E) Number of sampled time steps per worm, and (F) Sampling intervals for recorded neural activity. The horizontal dashed lines in (E) and (F) indicate the resampled sequence length ($L=314$) and time step ($\Delta t=0.333$ seconds) used in our preprocessing pipeline.
@@ 2.2 Calcium Fluorescence Data
We processed neural activity datasets from 11 open-source studies of C. elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [14, 5, 18, 37, 9, 15, 12, 10, 11, 13]. These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals. The number of worms and identified neurons varies across datasets as shown in Table 1. All worms were hermaphrodites at developmental stages no earlier than L4 [29], with the most being in adulthood (Figure 5).
@@
Index Dataset Name Num. Worms Mean Num. Neurons (labeled, recorded) Num. Labeled Neurons (min, max) Experimental Protocol
@@ 4 Kaplan2020 [9] 19 (36, 114) (23, 51) Wild-type; L4 stage; freely moving and immobilized
5 Yemini2021 [37] 49 (110, 125) (33, 179) Transgenic strain OH15500; adult stage; immobilized
6 Uzel2022 [18] 6 (50, 138) (46, 58) Well-fed wild-type; adult stage; immobilized
@@
8 Atanas2023 [5] 42 (88, 136) (64, 115) Transgenic strain NLS-GCaMP7f; adult stage; freely moving
9 Leifer2023 [14] 103 (69, 122) (26, 102) Wild-type; adult stage; immobilized
@@ 10 Lin2023 [11] 577 (8, 8) (1, 22) Wild-type; adult stage; immobilized
11 Venkatachalam2024 [32] 22 (187, 187) (185, 189) Wild-type; adult stage; freely moving
@@ 2.2.1 Overview of Neural Activity Datasets
We analyzed neural activity datasets collected from 11 independent studies, which varied in terms of the number of worms recorded, the number of neurons labeled, the duration of recordings, and sampling intervals. Figure 1 provides an overview of these datasets, highlighting their variability in experimental conditions.
@@ 2.2.2 Neural Data Processing
The neural activity data is preprocessed using a custom pipeline (Figure 2) which follows these steps:
1. Data Download and Extraction. The raw data from each dataset is downloaded from its respective source and extracted. The function download_url() downloads the dataset, and extract_zip() decompresses the data into a local directory while preserving the original folder structure.
2. Custom Dataset Implementations. Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format. The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking. Adding a new source dataset to the pipeline is as easy as defining a new child preprocessor class that implements any loading and processing logic that needs to be customized to that particular source; for example, if the new source’s data is stored in a special file type.
Figure 2: Processed Neural Activity Data Structure. Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C. elegans. The calcium signals are normalized, (optionally) smoothed, and resampled to a common time interval ($\Delta t$). A binary mask is generated to indicate which neurons were labeled for each worm.
We represent the neural activity data for a worm $k$ by a matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of time points and $D=300$ is the number of neurons in C. elegans. The neural activity at any time point $t$ is denoted as $\mathbf{X}^{(k)}[t]\in\mathbb{R}^{D}$, which represents the activity snapshot across the $D$ neurons.
3. Neuron Trace Normalization. The first preprocessing step is to normalize ($z$-score) the raw neural activity (i.e. data loaded from the source dataset files) across the neuron dimension. For each neuron trace $i$ in worm $k$, we compute:
$\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{raw}}-\mathbf{\mu}}{\mathbf{\sigma}}$
where $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$ are vectors of the mean and standard deviation, respectively, for each neuron, computed across the temporal dimension.
4. Smoothing. Optionally, the calcium traces can be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
$\mathbf{X}^{(k)}_{\text{smooth}}[t]=\alpha\mathbf{X}^{(k)}_{\text{norm}}[t]+(1-\alpha)\mathbf{X}^{(k)}_{\text{norm}}[t-1]$
where $\alpha$ is a smoothing hyperparameter set in preprocess.yaml. For our released neural dataset, smoothing was not applied ($\alpha=1$) as prior preprocessing had already been performed. Calcium fluorescence data inherently reflects low frequency neural dynamics. Smoothing, which acts further low pass over the raw data, is not needed.
5. Resampling. We resample the data to a fixed time interval $\Delta t$, using linear interpolation for datasets with different temporal resolutions. The new time index $t^{\prime}$ is calculated by:
@@
The resampled traces $\mathbf{X}_{\text{resampled}}^{(k)}$ are computed as:
$\mathbf{X}_{\text{resampled}}^{(k)}[t^{\prime}]=\mathbf{X}^{(k)}[t]+\left(\mathbf{X}^{(k)}[t+1]-\mathbf{X}^{(k)}[t]\right)\cdot(t^{\prime}-t)$
This ensures consistent temporal resolution across datasets. We use $\Delta t\approx 0.333$ seconds as a compromise to retain high- and low-frequency dynamics.
Importantly, the resampling step is the last preprocessing step necessary to create the neural activity data matrix for each worm:
@@
6. Masking and Subsetting. We also create a binary mask for each worm, $\mathbf{M}^{(k)}\in\{0,1\}^{D}$, where $D=300$ represents the total number of canonical neurons in C. elegans. Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled in worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not labeled.
@@
In the next section, we describe the second half the preprocessing pipeline, which unifies many measurements of the of C. elegans into a standardized format.
@@
for each dataset $\mathcal{D}_{\mathrm{raw,source}}$ in $\mathcal{D}_{\mathrm{raw,all\_sources}}$ do
     for each worm $k$ in $\mathcal{D}_{i}$ do
@@
         (Optional) Smooth traces: $\mathbf{X}^{(k)}_{\text{smooth}}\leftarrow\text{smooth}(\mathbf{X}^{(k)}_{\text{norm}},\alpha)$
         Resample traces: $\mathbf{X}^{(k)}_{\text{resample}}\leftarrow\text{resample}(\mathbf{X}^{(k)}_{\text{smooth}},\Delta t)$
@@
         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\texttt{boolean}(\mathbf{X}^{(k)}_{j}\neq\emptyset)\mid j\in[1,D]\}$
@@
     Save the processed dataset $\mathcal{D}_{\mathrm{processed,source}}:=\{\left(\mathbf{X}^{(k)},\mathbf{M}^{(k)}\right)\}_{k=1}^{N}$
@@ 2.3 Connectome Graph Data
We processed multiple open-source C. elegans connectome datasets, each capturing the neural wiring diagram, including both chemical synapses and electrical synapse (also referred as gap junctions). Table 5 provides metadata for these datasets, which vary in annotation precision, formats, and the extent of the nervous system covered.
@@ 2.3.1 Connectome Representation
  • Neuron Class and Type: Categories such as sensory, motor, or interneuron, and neuron-specific types (e.g., ADA, IL2, VD).
@@
  • Feature Vector ($\mathbf{x}_{i}$): Optionally, numerical data such as calcium fluorescence measurements over time.
@@
Edges are included in $\mathcal{E}$ if any of their attributes $\mathbf{a}_{ij}$ are nonzero. For example, an edge may represent gap junctions ($\mathbf{a}_{ij}=[0,g]$) or chemical synapses ($\mathbf{a}_{ij}=[c,0]$). In cases where functional connectivity is included, the graph may become fully connected ($|\mathcal{E}|\approx D^{2}$) unless significance thresholds are applied. Chemical and electrical synapse weights are based on counts from EM annotations [35, 3, 36] whereas the functional connectivity is based on signal propagation [14].

### v4

@@ Abstract
There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C. elegans), as this small model system provides a path to bridge the gap between nervous system structure (connectivity) and function (physiology). However, existing physiology datasets, whether involving passive recording or stimulation, are in distinct formats, and connectome datasets require preprocessing before analysis can commence. Here we compile and homogenize datasets of neural activity and connectivity. Our neural activity dataset is derived from 12 C. elegans neuroimaging experiments, while our connectivity dataset is compiled from 9 connectome annotations based on 3 primary electron microscopy studies and 1 signal propagation study. Physiology datasets, collected under varying protocols, measure calcium fluorescence in labeled subsets of the worm’s 300 neurons. Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling rate, yielding recordings from approximately 900 worms and 250 uniquely labeled neurons. The connectome datasets, collected from electron microscopy reconstructions, represent the entire nervous system as a graph of connections. Our collection is accessible on HuggingFace, facilitating analysis of the structure-function relationship in biology using modern neural network architectures and enabling cross-lab and cross-animal comparisons.
@@ 1 Introduction
To address these challenges, we have created two integrated datasets. The first aggregates neural dynamics from 12 calcium imaging studies, and the second compiles connectome data from three electron microscopy wiring studies [37, 4, 38] and one functional signal propagation study [15]. Together, these datasets combine structure and function, offering an invaluable resource for computational modelers aiming to construct biologically grounded neural network models.
@@ 2.1 Code Repository
We provide an open-source code repository that implements the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions, files and links) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data into standardized formats.
@@
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The pickle_neural_data() function extracts and processes raw calcium fluorescence data from each of our 12 neural activity source datasets, standardizing them into neural activity matrices that are saved as compressed files for efficient storage and retrieval. The preprocess_connectome() function processes raw connectome data from source files into graph tensors, which are formatted for compatibility with graph neural network (GNN) frameworks such as PyTorch Geometric [22]. This preprocessing step facilitates downstream tasks like connectivity-informed modeling of neural activity. The 10 distinct connectome source files are detailed in Table 5.
Our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
Figure 1: Overview of Neural Activity Datasets. (A) Distribution of the number of worms in each experimental dataset. (B) Number of recorded neurons per worm compared to the total neuronal population of C. elegans. (C) Total duration of recorded neural activity for each dataset. (D) Average recording duration per worm, with one hour of calcium imaging as a benchmark. (E) Number of resampled time steps per worm, and (F) Pre-resampled sampling intervals for recorded neural activity. The horizontal dashed line in (F) indicates the target resampled time step ($\Delta t=0.333$ seconds) used in our preprocessing pipeline.
@@ 2.2 Calcium Fluorescence Data
We processed neural activity datasets from 12 open-source studies of C. elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [15, 6, 19, 39, 10, 16, 13, 11, 12, 14]. These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals. The number of worms and identified neurons varied across datasets as shown in Table 1. All worms were hermaphrodites at developmental stages no earlier than L4 [31], with most being in adulthood (Figure 5).
@@
Index Dataset Name Num. Worms
Mean Num.
Neurons
(labeled, recorded)
Num. Labeled
Neurons
(min, max)
Experimental Protocol
@@ 4 Kaplan2020 [10] 19 (36, 114) (23, 51) Wild-type; L4 stage; freely moving and immobilized
5 Nejatbakhsh2020 [29] 21 (173, 175) (163, 184) Transgenic (GCaMP6); L4 to adult stages; during chemotaxis/decision-making tasks; immobilized
6 Yemini2021 [39] 49 (110, 110) (33, 179) Transgenic (OH15500); adult stage; immobilized
7 Uzel2022 [19] 6 (50, 138) (46, 58) Wild-type; adult stage; Well-fed but immobilized
@@
9 Leifer2023 [15] 110 (63, 67) (13, 98) Wild-type; adult stage; immobilized
@@ 10 Lin2023 [12] 577 (8, 8) (1, 22) Wild-type; adult stage; immobilized
11 Flavell2023 [1] 40 (88, 136) (64, 115) Transgenic (GCaMP7f, mNeptune2.5); adult stage; freely moving active and quiescent states
12 Venkatachalam2024 [34] 22 (187, 187) (185, 189) Wild-type; adult stage; immobilized in microfluidic chips
@@ 2.2.1 Overview of Neural Activity Datasets
We analyzed neural activity datasets collected from 12 independent studies, which varied in terms of the number of worms recorded, the number of neurons labeled, the duration of recordings, and sampling intervals. Figure 1 provides some summary statistics on the neural activity data collected from each of these datasets.
2.2.2 Brief Note on Notation
In this section, we define the key mathematical notation used throughout our preprocessing pipeline and neural activity data representation. This note aims to provide clarity and consistency for readers, particularly when interpreting symbols used in equations and methods. For a comprehensive description of the symbols, including those specific to other parts of the pipeline, refer to Appendix Table 5.
We represent the neural activity data for a worm $k$ as a time-series matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of recorded time points, and $D=300$ is the total number of canonical neurons in C. elegans. This excludes the canal-associated neurons (CAN), which are not typically included in functional studies [18, 33]. Each row of $\mathbf{X}^{(k)}$ corresponds to a snapshot of neural activity across all $D$ neurons at a single time point, while each column corresponds to the activity trace for a specific neuron over time.
Key conventions include:
  • Activity Snapshot: The neural activity at time $t$ is denoted $\mathbf{X}^{(k)}[t]\in\mathbb{R}^{D}$, representing the activity of all $D$ neurons at that moment.
  • Indexing: We use 00-based indexing for time points (e.g., $t=0,1,\dots,T_{k}-1$) and $1$-based indexing for neurons (e.g., $i=1,2,\dots,D$).
  • Vector Orientation: All vectors (e.g., $\mathbf{X}^{(k)}[t]$) are column vectors unless explicitly stated otherwise.
This notation provides the foundation for the mathematical operations described in the neural data processing pipeline. For additional details on notation, see Appendix Table 5.
@@ 2.2.3 Neural Data Processing
The neural activity data is preprocessed using a custom pipeline (Figure 2) which follows the general outline given in Algorithm 1 and described in more detail next.
a). Data Download and Extraction. The raw data from each dataset is downloaded from its respective source and extracted. The function download_url() downloads the dataset, and extract_zip() decompresses the data into a local directory while preserving the original folder structure.
b). Custom Dataset Implementations. Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format. The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps in order: resampling, optional smoothing, and normalization. Adding a new source dataset to the pipeline is as easy as defining a new child preprocessor class that implements any loading and processing logic that needs to be customized to that particular source; for example, if the new source’s data is stored in a special file type.
Figure 2: Processed Neural Activity Data Structure. Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C. elegans. The calcium signals are resampled, (optionally) smoothed, and normalized to the same rate ($1/\Delta t$ Hz). A binary mask is generated to indicate which neurons were labeled for each worm.
c). Resampling. The first preprocessing step resamples the raw data to a fixed time interval $\Delta t$, using linear interpolation for datasets with different temporal resolutions. The new time index $t^{\prime}$ is calculated by:
@@
The resampled traces $\mathbf{X}_{\text{resampled}}^{(k)}$ are computed from the raw data $\mathbf{X}^{(k)}_{\text{raw}}$ as:
$\mathbf{X}_{\text{resampled}}^{(k)}[t^{\prime}]=\mathbf{X}^{(k)}_{\text{raw}}[t]+\left(\mathbf{X}^{(k)}_{\text{raw}}[t+1]-\mathbf{X}^{(k)}_{\text{raw}}[t]\right)\cdot(t^{\prime}-t)$
This ensures consistent temporal resolution across datasets. We use $\Delta t\approx 0.333$ seconds as a compromise to retain high- and low-frequency dynamics. After resampling, we set:
@@
Resampling is done as the first preprocessing step to ensure meaningful cumulative data, as we’ll discuss later on in step 5 of preprocessing, normalization.
d). Optional Smoothing. Following resampling, the calcium traces can optionally be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
$\mathbf{X}^{(k)}_{\text{smooth}}[t]=\alpha\mathbf{X}^{(k)}[t]+(1-\alpha)\mathbf{X}^{(k)}[t-1]$
where $\alpha$ is a smoothing hyperparameter set in preprocess/config.py. For our released neural dataset, smoothing was not applied ($\alpha=1$) as prior preprocessing had already been performed. Calcium fluorescence data inherently reflects low frequency neural dynamics. Smoothing acts as a further low pass over the resampled data and may therefore be unnecessary.
e). Neuron Trace Normalization. As the final preprocessing step, the preprocessors normalize (or $z$-score) the neural activity across the neuron dimension. Let $\mathbf{X}^{(k)}_{\text{input}}:=\mathbf{X}^{(k)}_{\text{smooth}}$ if smoothing was applied; otherwise $\mathbf{X}^{(k)}_{\text{input}}:=\mathbf{X}^{(k)}_{\text{resampled}}$. For each neuron trace $i$ in worm $k$, we compute:
$\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{input}}-\boldsymbol{\mu}}{\boldsymbol{\sigma}},$ (1)
where $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$ are column vectors of the mean and standard deviation, respectively, for each neuron, calculated across the temporal dimension.
This standard normalization method uses the full time-series to compute $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$, which introduces information from future states into earlier time points, thereby violating causality. While this may be acceptable for some applications, respecting causality could be crucial in certain cases, such as for real-time or streaming neural activity models.
f). Masking and Subsetting. We also create a binary mask for each worm, $\mathbf{M}^{(k)}\in\{0,1\}^{D}$, where $D=300$ represents the total number of canonical neurons in C. elegans. Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled in worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not labeled.
@@
for each dataset $\mathcal{D}_{\mathrm{raw,source}~{}i}$ in $\mathcal{D}_{\mathrm{raw,all\_sources}}$ do
     for each worm $k$ in $\mathcal{D}_{\mathrm{raw,source}i}$ do
@@
         if Smooth traces then
              $\mathbf{X}^{(k)}_{\text{original}}\leftarrow\text{smooth}(\mathbf{X}^{(k)}_{\text{norm}},\alpha)$
         else
              $\mathbf{X}^{(k)}_{\text{original}}\leftarrow\mathbf{X}^{(k)}_{\text{norm}}$
         end if
         Resample traces: $\mathbf{X}^{(k)}_{\text{resample}}\leftarrow\text{resample}(\mathbf{X}^{(k)}_{\text{original}},\Delta t)$
@@
         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\texttt{boolean}(\mathbf{X}^{(k)}_{j}\neq\boldsymbol{\emptyset})\mid j\in[1,D]\}$
@@
     Save the processed dataset $\mathcal{D}_{\mathrm{processed,source}~{}i}:=\{\left(\mathbf{X}^{(k)},\mathbf{M}^{(k)}\right)\}_{k=1}^{N}$
@@ 2.3 Connectome Graph Data
We processed multiple open-source C. elegans connectome datasets, each capturing the neural wiring diagram, including both chemical synapses and electrical synapse (also known as gap junctions). Table 5 provides metadata for these datasets, which vary in annotation precision, formats, and the extent of the nervous system covered.
@@ 2.3.1 Connectome Representation
  • Neuron Class and Type: Categories such as sensory, motor, or interneuron, and neuron-specific types (e.g., ADA, IL2, VD1).
@@
  • Feature Vector ($\mathbf{x}_{i}$): Optionally, numerical data (e.g. calcium fluorescence measurements over a fixed time window, gene expression embeddings)
@@
Edges are included in $\mathcal{E}$ if any of their attributes $\mathbf{a}_{ij}$ are nonzero. For example, an edge may represent gap junctions ($\mathbf{a}_{ij}=[0,g]$) or chemical synapses ($\mathbf{a}_{ij}=[c,0]$). In cases where functional connectivity is included, the graph may become fully connected ($|\mathcal{E}|\approx D^{2}$) unless a significance threshold is applied to filter out weak functional connections. Chemical and electrical synapse weights are based on counts from EM annotations [37, 4, 38] whereas the functional connectivity weights are based on signal propagation [15].